OpenStudy (anonymous):

PLEASE HELP FAST! How do i find the width? 3w + L = 56 ft?!

OpenStudy (the_fizicx99):

W = 56-L/3

OpenStudy (the_fizicx99):

Isolating it is what you meant, right?

OpenStudy (anonymous):

yes and no, i need to know like the actual answer. in all im trying to figure out what the length is. but it says i have to find out what the width is first

OpenStudy (the_fizicx99):

Well the width would be 1866 --> 18.7

OpenStudy (the_fizicx99):

18.66*

OpenStudy (anonymous):

okay. so...from there how would i find out the length?

OpenStudy (the_fizicx99):

Hm lol thats where I'm stuck

OpenStudy (anonymous):

i know same D;

OpenStudy (the_fizicx99):

This is grade..?

OpenStudy (anonymous):

well maybe 56 - 18.7? 37.3?

OpenStudy (anonymous):

yes its for a grade

OpenStudy (the_fizicx99):

lol what grade, what're the options

OpenStudy (anonymous):

what? aha what do you mean? this is an assignment question. i am trying to find the length. thats all i need to find out, i dont have any options

OpenStudy (the_fizicx99):

ask @phi

OpenStudy (phi):

where did you get the equation from How do i find the width? 3w + L = 56 ft ?

OpenStudy (anonymous):

the perimeter is 56. the width is 3 times the length. L=3w. 3w + L = 56 ft

OpenStudy (phi):

You should say perimeter of a rectangle is 2(L+W) which, in this case, equals 56 2(L+W)=56 you say ***the width is 3 times the length*** (which means W= 3L) and you say ***L=3w (which means length is 3 times the width which is it? let's say W= 3L replace W with 3L in the first equation 2(L+3L)= 56 now solve for L

OpenStudy (the_fizicx99):

Ohh, missed that